Use the thermodynamic data provided below to estimate the boiling point (in K) of V. Report your answer to zero decimal places in standard notation (i.e. 123. kJ *For numbers ending in zero, be sure to include the decimal!*).

Substance ΔH°f (kJ/mol) S° (J mol-1K-1)
V (l) 17.3 36.1
V (g) 515.5 182.3
Answer #1

Step 1:

Given:

Hof(V(l)) = 17.3 KJ/mol

Hof(V(g)) = 515.5 KJ/mol

Balanced chemical equation is:

V(l) ---> V(g)

ΔHo rxn = 1*Hof(V(g)) - 1*Hof( V(l))

ΔHo rxn = 1*(515.5) - 1*(17.3)

ΔHo rxn = 498.2 KJ

Step 2:

Given:

Sof(V(l)) = 36.1 J/mol.K

Sof(V(g)) = 182.3 J/mol.K

Balanced chemical equation is:

V(l) ---> V(g)

ΔSo rxn = 1*Sof(V(g)) - 1*Sof( V(l))

ΔSo rxn = 1*(182.3) - 1*(36.1)

ΔSo rxn = 146.2 J/K

Step 3:

Boiling is an equilibrium process and hence ΔGo = 0

ΔGo = 0.0 KJ/mol

ΔHo = 498.2 KJ/mol

ΔSo = 146.2 J/mol.K

= 0.1462 KJ/mol.K

use:

ΔGo = ΔHo - T*ΔSo

0.0 = 498.2 - T *0.1462

T = 3408. K

Answer: 3408 K